# Oracle-Based Multistep Strategy for Solving Polynomial Systems Over Finite Fields and Algebraic Cryptanalysis of the Aradi Cipher

### v1

@@ 7. Conclusions and further directions
In this paper, we propose a new version of the multistep solving strategy, based on a recursive Depth-First Search implementation and the introduction of an oracle function designed to predict whether a polynomial system requires further simplification through variable evaluations. This approach enables a simple yet effective implementation of the MultiSolve algorithm and supports a complexity analysis based on tree structures. Furthermore, it provides a unified framework that generalizes various solving strategies, including the one used in our attack on the Trivium cipher and the classic hybrid approach, into a single algorithm, where only the oracle function needs to be adapted.
Thanks to this work, we were able to carry out an algebraic attack on the recent block cipher Aradi revealing some critical vulnerabilities. Notably, our attack is the first to target all 16 rounds of the cipher.
As a further direction for this line of research, we suggest developing oracle functions based on Machine Learning or Deep Learning algorithms, trained on the ideals, and corresponding Macaulay matrices, generated by MultiSolve. A preliminary experimental study is currently underway. Another promising direction is the development of a parallel implementation of the MultiSolve algorithm for High Performance Computing environments, which could significantly improve practical efficiency and broaden the applicability of the multistep strategy to the algebraic cryptanalysis of other relevant cryptographic systems.

### v3

@@ 7. Conclusions and further directions
In this paper, we propose a new version of the multistep solving strategy, based on a recursive Depth-First Search implementation and on the introduction of an “oracle function” intended to predict whether a polynomial system requires further simplification through variable evaluations.
This approach enables a simple yet effective implementation of the MultiSolve algorithm and supports a complexity analysis based on tree structures. Furthermore, it provides a unified framework that generalizes various solving strategies, including the one used in our attack on the Trivium cipher and the classic hybrid approach, into a single algorithm, where only the oracle function needs to be adapted. Thanks to this work, we were able to perform an algebraic attack on the recent block cipher Aradi, which revealed some aspects of its symbolic modeling.
As a further direction for this line of research, we suggest developing oracle functions based on Machine Learning or Deep Learning algorithms, trained on the ideals, and corresponding Macaulay matrices, generated by MultiSolve. A preliminary experimental study is currently underway. Another promising direction lies in the development of a parallel implementation of the MultiSolve algorithm, as well as in its extension to polynomial systems admitting multiple solutions, which could be leveraged in discrete optimization tasks.